In a friendly game of handball, you hit the ball essentially at ground level and send it toward the wall with a speed of 26 m/s at an angle of 34 ∘ above the horizontal.

a)

How long does it take for the ball to reach the wall if it is 3.3 m away?

b)

How high is the ball when it hits the wall?
Answer #1

Here,V =26m/s

theta=34

Vx = Vcos34 = 21.55m/s

Horizontal (x) distance from wall =3.3 m

Time taken = horizontal distance/Vx = 0.153 s

Part B

Vy = Vsin34 = 14.54m/s

using formula, Sy= Vyt+1/2 at​​​​​2​​​​at2

= 14.54×0.153m - 9.8×(0.153)2 /2

= 2.225m -0.115 m = 2.11 m